Identifying and Analyzing Financial Statement Effects of Dividends
The stockholders' equity of Kinney Company at December 31, 2011, is shown below.

5% preferred stock, $100 par value, 10,000 shares authorized; 4,000 shares issued and outstanding $ 400,000
Common stock, $5 par value, 200,000 shares authorized; 50,000 shares issued and outstanding 250,000
Paid-in capital in excess of par value—preferred stock 40,000
Paid-in capital in excess of par value—common stock 300,000
Retained earnings 656,000
Total stockholders' equity $1,646,000


The following transactions, among others, occurred during 2012:
Apr. 1 Declared and issued a 100% stock dividend on all outstanding shares of common stock. The market value of the stock was $11 per share.
Dec. 7 Declared and issued a 3% stock dividend on all outstanding shares of common stock. The market value of the stock was $14 per share.
Dec. 20 Declared and paid (1) the annual cash dividend on the preferred stock and (2) a cash dividend of 80 cents per common share.

(a) Use the financial statement effects template to indicate the effects of these separate transactions.

Use negative signs with answers, when appropriate.

Balance Sheet

Transaction Cash Asset +

Noncash

Assets

= Liabilities +

Contributed

Capital

+

Earned

Capital

Apr. 1 Answer Answer Answer Answer Answer
Dec. 7 Answer Answer Answer Answer Answer
Dec. 20 Answer Answer Answer Answer Answer

Income Statement


Revenue

-

Expenses

=

Net

Income

Answer Answer Answer
Answer Answer Answer
Answer Answer Answer


(b) Compute retained earnings for 2012 assuming that the company reports 2012 net income of $253,000.
$
Answer #1

(a) Use the financial statement effects template to indicate the effects of these separate transactions.

Use negative signs with answers, when appropriate.

Balance Sheet

Transaction Cash Asset +

Noncash

Assets

= Liabilities +

Contributed

Capital

+

Earned

Capital

Apr. 1 0 + 0 = 0 + 250000 + (250000)
Dec. 7 0 + 0 = 0 + 42000 + (42000)
Dec. 20 (102400) + 0 = 0 + 0 + (102400)

April 1= $250000*100%= $250000

Stock dividend shares on April 1= 50000*100%= 50000

Stock dividend shares on Dec 7= (50000+50000*100%)*3%= 3000

Stock dividend= 3000*$14= $42000

Preferred dividend= $400000*5%= $20000

Total outstanding common shares= Beginning shares+Stock dividend shares on April 1+Stock dividend shares on Dec 7

= 50000+50000+3000= 103000 shares

Common stock dividend= 103000*80%= $82400

Total dividend= Preferred dividend+Common stock dividend= $20000+82400= $102400

Income Statement


Revenue

-

Expenses

=

Net

Income

0 0 0
0 0 0
0 0 0


(b) Compute retained earnings for 2012 assuming that the company reports 2012 net income of $253,000

Beginning balance of Retained earnings $656000
Add: Net income 253000
909000
Less: Stock dividend (250000+42000) 292000
Less: Cash dividend (20000+82400) 102400 (394400)
Ending balance of Retained earnings $514600